OpenStudy (anonymous):

y=4x how can I make 3 ordered pairs?

OpenStudy (katrinakaif):

Its been a while, But i think you can just put in any x and y coordinate but your slope has to say the same. To find the equation use the formula (y-y1) = m(x-x1) as x1 and y1 are the coordinates you picked

OpenStudy (anonymous):

I dont understand that katrina m(x-x1)?

OpenStudy (katrinakaif):

Okay so lets say we picked random coordinates of (1,1) also equal to be (x1, y1) You have teh slope of 4 which is m in the following equation (y-y1) = m(x-x1) ==> (y-1) = 4(x-1) Distribute and solve until you get a form of y = mx + b You got your equation

OpenStudy (anonymous):

thanx